### 08 Apr 2024 v2

# Coherent Perfect Absorption of Arbitrary Wavefronts at an Exceptional Point

Abstract
A Coherent Perfect Absorber (CPA) exploits the interferometric nature of light to deposit all of a light field’s incident energy into an otherwise weakly absorbing sample. The downside of this concept is that the necessary destructive interference in CPAs gets easily destroyed both by spectrally or spatially detuning the incoming light field. Each of these two limitations has recently been overcome by insights from exceptional-point physics and by using a degenerate cavity, respectively. Here, we show how these two concepts can be combined into a new type of cavity design, which allows broadband exceptional-point absorption of arbitrary wavefronts. We present two possible implementations of such a Massively Degenerate Exceptional-Point absorber and compare analytical results with numerical simulations.
preprint: APS/123-QED
Figure 1: (a) A conventional CPA (critically coupled cavity). When the critical-coupling condition $R_{1}=T_{\mathrm{a}}^{2}$ is met, the incident coherent light-field is perfectly absorbed, but only for a single plane-wave input mode at the right frequency. (b) A spectrally degenerate EP-CPA, consisting of two weakly coupled conventional CPAs. When the critical-coupling-conditions $R_{\mathrm{c}}=4R_{1}/\left(1+R_{1}\right)^{2}$ and $R_{1}=T_{\mathrm{a}}^{2}$ are met, the (generally distinct) resonance-points merge at a single, real frequency. This spectral degeneracy results in a broadened (quartic) absorption spectrum (not shown). (c) A MAD-CPA is a spatially-degenerate extension of the conventional CPA: Because of the self-imaging telescopic lens-arrangement, any incoming spatial mode is mapped onto itself after each round-trip, leading to perfect cancellation of any back-reflected light at the front mirror. (d) The MAD-EP-CPA (“Design A”) combines the features of spatial and spectral degeneracies for improved spatial and spectral acceptance.
Introduction. — While thin films or layers can only absorb a certain percentage of the light that shines through them in a single pass, an enhancement of absorption can be reached when the light beam has multiple passes through the absorbing material. This insight already indicates that placing an absorber in a cavity, where light bounces back and forth, can enhance the degree of absorption considerably (see Fig. 1(a) as an example). The interferometric nature of light propagation in a cavity can boost the absorption even further – up to the point, where all the incoming light gets perfectly absorbed. In the single-mode case this phenomenon is known as “critical coupling” [1, 2, 3, 4, 5]. For multiple modes one speaks of “coherent perfect absorption”, an effect that can also be interferometrically controlled through the relative phases of the incoming modes [6, 7, 8, 9, 10, 11, 12, 13]. CPAs have been widely explored in diverse platforms and applications, encompassing complex structures and disordered media [14, 15, 16, 17, 18], single-port interferometers [19, 20], optical switches [21, 22], sensors [23, 24], all-optical transistors and logic gates [25, 26]. Recently, two of the major challenges of coherent perfect absorbers (CPAs) have been successfully addressed:
(i) The first challenge is a CPA’s narrow bandwidth, with slight frequency shifts disrupting the resonant absorption, typically resulting in a Lorentzian absorption profile. Research on how to make CPAs more broadband encompasses a wide array of approaches, such as metasurfaces [27, 28, 29], plasmonic effects [30, 31, 32], metallic nanoparticles metasurfaces [33], thin films [34], and broadband PCN resonators [35]. Recently also insights from Exceptional Point (EP) physics have been used for advancing this goal [11, 12, 13, 36]: At an EP, two (or more) discrete eigenvalues and eigenstates of a non-Hermitian system merge. Such an EP-CPA can be observed when two (or more) cavities with spectrally-overlapping resonances are critically coupled as depicted in Fig. 1(b). Specifically, it is possible to select system parameters such that both the critical-coupling condition, neccessary for perfect absorption, and the EP-condition coincide at the same real input frequency [11]. The result is a spectral degeneracy that leads to a significantly broadened (quartic) absorption spectrum [12, 13], in analogy to the physics of white-light cavities [37, 38, 39, 40, 41, 42] and of multi-mirror Fabry-Pérot interferometers [43].
(ii) The second major challenge concerns the limitation that a CPA typically only works for a well-defined input wavefront or mode. Detuning this input mode, e.g., in its incoming angle or phase profile, destroys the delicate interference necessary to achieve perfect absorption. To overcome this restriction, a massively degenerate (MAD) cavity turns out to be ideally suited [44], as illustrated in Fig. 1(c). In such a cavity, every incident light-field is always imaged onto itself after each round-trip. This configuration thus enables the necessary interferometric CPA-effect for any given spatial mode.
The question we explore here is whether these two distinct concepts can be integrated into a single CPA design that simultaneously achieves spectral and spatial degeneracy. Our approach to this new cavity design is based on the following considerations: A conventional CPA, as depicted in Fig. 1(a), can be transformed into a spectrally degenerate CPA by critically coupling it to another cavity, as illustrated in Fig. 1(b). Similarly, the same conventional CPA can be transformed into a spatially degenerate CPA by incorporating two lenses (focal length $f$) in a telescopic configuration (of length $4f$), as illustrated in Fig. 1(c). Thus, merging these two ideas by critically coupling two spatially degenerate cavities, as illustrated in Fig. 1(d), should result in a CPA that is both spatially and spectrally degenerate.
Spatial and Spectral Degeneracy. — To derive the reflection behavior of light incident from the left on this MAD-EP-CPA we employ a scalar optics model for polarized light, considering a large, yet finite, number of spatial modes. Let $\textbf{R}_{1}$, $\textbf{R}_{2}$, and $\textbf{R}_{\mathrm{c}}$ represent the reflection matrices, and $\textbf{T}_{1}$, $\textbf{T}_{2}$, and $\textbf{T}_{\mathrm{c}}$ the transmission matrices for the left, right and central mirror, respectively. Additionally, $\textbf{T}_{\mathrm{a}}$ is the transmission through the absorber and $\textbf{T}_{\text{4f}}$ represents the transmission matrix for propagation through a $4f$-system, comprising two lenses in a telescopic configuration (without mirrors or absorbers). $\textbf{T}_{\text{4f}}$ acts as a double Fourier transform producing a spatially flipped image and a phase shift $e^{i\phi}$ that is the same for all modes. Using that the same reflection and transmission coefficients apply to all modes at the mirrors and the absorber, all matrices except $\textbf{T}_{\text{4f}}$ become diagonal matrices, e.g., $\textbf{R}_{1}=r_{1}\mathbb{1}$, $\textbf{T}_{1}=t_{1}\mathbb{1}$, etc. Using a transfer matrix approach, we can construct from these individual transmission matrices the sought-after reflection matrix $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(A)}}}$ for the entire MAD-EP-CPA [45, S1] (Design A in Fig. 1(d)):
$\textbf{R}_{\text{CPA}}^{\text{(A)}}=\frac{t_{\text{a}}^{2}\,\textbf{T}_{\text{4f}}^{4}+r_{\text{c}}(1+r_{1}t_{\text{a}}^{2})\,\textbf{T}_{\text{4f}}^{2}+r_{1}\mathbb{1}}{r_{1}t_{\text{a}}^{2}\,\textbf{T}_{\text{4f}}^{4}+r_{\text{c}}(r_{1}+t_{\text{a}}^{2})\,\textbf{T}_{\text{4f}}^{2}+\mathbb{1}}$ (1)
The matrix $\textbf{T}_{\text{4f}}^{2}$ in Eq. (1) results in a quadruple Fourier transformation with a uniform phase shift $e^{2i\phi}$ for all modes and can thus be represented as a simple diagonal matrix $\textbf{T}_{\text{4f}}^{2}=e^{2i\phi}\mathbb{1}$. Analogously, $\textbf{T}_{\text{4f}}^{4}=e^{4i\phi}\mathbb{1}$, turning $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(A)}}}$ into a diagonal matrix with identical entries $r_{\text{CPA}}(\omega)$ along the diagonal. Therefore, Eq. (1) can be reduced to a simple scalar equation, which proves that all modes are subject to the same reflection behavior in the system (spatial degeneracy). The function $r_{\text{CPA}}(\omega)$ is determined by the power-reflectivities $R_{1}=|r_{1}|^{2}$, $R_{\text{c}}=|r_{\text{c}}|^{2}$ of the mirrors, and by the absorber’s one-way transmissivity $T_{\text{a}}=|t_{\text{a}}|^{2}$. To determine the conditions for spectral degeneracy, i.e. for the EP to occur at a real frequency, the values $R_{1}$, $R_{\mathrm{c}}$ and $T_{\mathrm{a}}$ need to be adjusted such that the two solutions $\omega_{1}$, $\omega_{2}$ of the equation $r_{\mathrm{CPA}}(\omega)=0$ merge into a single real-valued frequency $\omega$. These conditions are:
$R_{1}=T_{\mathrm{a}}^{2}\;\;\;\text{and}\;\;\;R_{\mathrm{c}}=\frac{4R_{1}}{\left(1+R_{1}\right)^{2}}$ (2)
A detailed derivation is available in the supplementary material [45, S1-S3]. The result (2) shows that the relationship between $T_{\mathrm{a}}$ and $R_{1}$ for a MAD-EP-CPA is identical to that of a single-cavity MAD-CPA [44]: The higher the reflectivity $R_{1}$ of the input-coupling mirror, the weaker the internal absorber needs to be. Additionally, the relation between $R_{\text{c}}$ and $R_{1}$ indictates that $R_{\text{c}}$ quickly approaches values near $1$ with increasing values of $R_{1}$, indicating weak coupling between the left and right sub-cavity.
Numerical Simulations. — To corroborate that these analytical results are, indeed, applicable to the multi-mode MAD-EP-CPA, and to assess the impact of deviations in certain system parameters (such as minor misalignments of the mirrors), we conducted a corresponding numerical Scalar Fourier Optics computer simulation. As input we used a random speckle field, see Fig. 2(a), generated by 160 modes of equal amplitude and random phase in $k$-space. Without compensating for the refractive index of the absorber (which we assumed to be $n_{r}=1.5$, at a thickness of $d=0.6\mathrm{mm}$), our numerical simulation already indicates a considerably broadened absorption lineshape, but with the minimum reaching only 0.7%, see the red line in Fig. 2(b). To address this, we adjusted the focal length of the rightmost lens in Fig. 1(d) to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$. This adjustment compensates for the increased optical path-length caused by the absorber and preserves identical optical path-lengths in both the left and right sub-cavity. As shown in the numerically calculated reflection spectrum of Fig. 2(b) (blue line), this correction restores perfect absorption at the point of degeneracy and verifies the successful operation of our MAD-EP-CPA cavity design (see the broad yellow line in Fig. 2(b) for comparison with an ideal quartic absorption lineshape). For a more detailed explanation of this refraction compensation technique, readers are referred to the supplementary text [45, S8]. Another potential strategy to minimize refractive aberrations is to use an absorber with a refractive index whose real part is as close to $1$ as possible [46].
Figure 2: Numerical simulation of the proposed EP-MAD-CPA, as depicted in Fig. 1(d). (a) Intensity pattern of a simulated complex input random speckle field, composed of $160$ modes. (b) Reflection spectrum across the entire field-of-view without compensation of the refraction aberrations caused by the absorber with thickness $0.6\mathrm{mm}$ (red line), and with compensation of this shift included (blue line). For reference, ideal quartic behavior is shown with a thick yellow line, and a conventional CPA with the same absorber with a black dotted line. (c) Reflected field intensity pattern at the resonance point when the refraction in the absorber is not compensated. (d) Reflected field intensity pattern at the resonance point with refraction compensated. (Simulation parameters: $R_{1}=0.7$, $R_{\mathrm{c}}=0.96886$. $T_{\mathrm{a}}=\sqrt{0.7}$, wavelength $\lambda_{0}=633\mathrm{nm}$, focal length $f=75\mathrm{mm}$.)
The numerical simulations also allow us to study the impact of small deviations from optimal system parameters, and help to pinpoint the critical system parameters influencing the system performance. For instance, the simulations demonstrate that the system exhibits robustness to deviations of the optimal reflectivity $R_{1}$ of the input coupling mirror (to within about ±$\pm 5\%$), or of the optimal transmissivity $T_{\mathrm{a}}$ of the absorber (to within about ±$\pm 2\%$). In contrast, deviations in the reflectivity $R_{\mathrm{c}}$ of the central coupling mirror from its optimal value significantly affect the system performance. Even a mere $1\%$ deviation from the optimal value results in a minimum reflectivity of approximately $5\%$, significantly diminishing the system’s absorption efficiency (further details and figures can be found in the supplementary text [45, S4]). As expected, the precise parallel alignment of the mirrors is also crucial, particularly for the central coupling mirror, as shown in Fig. 3. Additional details on critical parameters, such as the consequences of slight discrepancies in the alignment of the two cavity lengths or of residual lens reflections, are available in the supplementary text [45, S4].
We have also studied MAD-EP-CPA systems encompassing more than two coupled cavities, as detailed in the supplementary text [45, S5]. Interestingly, constructing an EP-CPA with an increasing number of cavities results in a reflection spectrum around the resonance frequencies akin to a Butterworth bandpass filter, a concept from electrical filter theory [47, 48]. This observation is consistent with the established understanding that, as outlined in [43], the Butterworth function also characterizes the transmission behavior of a multi-mirror Fabry-Pérot interferometer in the absence of a dissipative element.
Figure 3: Sensitivity to misalignment of the central coupling mirror (same simulation parameters as in Fig. 2). Left: Reflection spectrum across the entire field-of-view when the center mirror is slightly tilted out of the vertical alignment. Right: Reflected field intensity pattern when the central mirror is tilted by $0.001^{\circ}$. Perfect absorption now only occurs in a narrow horizontal area along the tilt axis.
Alternative implementation. — The initial design for a MAD-EP-CPA (“Design A”), as discussed thus far and illustrated in Fig. 1(d), is probably the most obvious solution for merging the concepts shown in Fig. 1(b) and (c).
Figure 4: Alternative MAD-EP-CPA configuration (“Design B”). (a) Rather than coupling two $4f$-cavities as in Design A, a coupling mirror ($R_{\mathrm{c}}$) is positioned at the center of a single $4f$-cavity. The left mirror is divided into an input-coupling-part ($R_{1\mathrm{a}}$) illustrated above the optical axis, and a completely reflecting mirror ($R_{1\mathrm{b}}$) depicted below the optical axis. The absorber with transmissivity $T_{\mathrm{a}}$ is placed next to mirror $R_{1\mathrm{b}}$. To correct for refraction aberrations, the left lens’s focal length is adjusted to $f^{\prime}$, and a transparent slab $T_{1}$, matching $T_{\text{a}}$ in thickness and real value of refractive index, is placed near $R_{1\mathrm{a}}$. Exemplary beam paths are marked in red and blue above and below the optical axis. (b) Numerical simulation of the field-of-view reflection spectrum (blue line) in comparison to an ideal quartic behavior (broad yellow line). (c) Input field intensity pattern. (d) Reflected field intensity pattern, demonstrating perfect absorption.
However, Design A poses challenges in the implementation, as it involves weakly coupling two plane-parallel cavities with altogether four embedded lenses. We thus propose in Fig. 4(a) an alternative “Design B”, involving a coupling mirror with reflectivity $R_{\mathrm{c}}$ at the center of a single $4f$-cavity, requiring only two lenses. This simplification requires a division of the left mirror at the optical axis into an input-coupling segment with reflectivity $R_{1\mathrm{a}}<1$ and a fully reflective mirror $R_{1\mathrm{b}}=1$. The weak absorber is then placed adjacent to the totally reflective mirror $R_{1\mathrm{b}}$. Additionally, to mitigate the additional refraction introduced by the absorber, we adjust the focal length of the left lens to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$ (see supplementary [45, S8]). Furthermore, to ensure an identical optical path-length above and below the optical axis, we have incorporated a transparent slab, designated $T_{1}$, having the same thickness and refractive index as the absorber, adjacent to mirror $R_{1\mathrm{b}}$.
Design B can be seen as a “folded-up” variant of Design A by identifying analogous round-trip paths. For instance, in Design A, light undergoes self-imaging round-trips within the right sub-cavity, with a corresponding path in Design B from the bottom-left to the top-right. A similar parallel can be observed for round-trips in Design A’s left sub-cavity, corresponding with the path from top-left to bottom-right in Design B. In both of these scenarios, a single reflection at Design A’s central mirror is substituted by two transmissions through Design B’s central mirror. Also, for the third round-trip-path in Design A, where light propagates between the farthest mirrors $R_{1}$ and $R_{2}$, an analog round-trip-path can be identified in Design B (see supplementary [45, S7] for details). However, it is crucial to recognize that such comparisons between Design A and Design B offer, at best, a qualitative analogy. A comprehensive comparison requires a sound mathematical analysis, considering the infinite sum of all possible round trips and the effects of coupling between the two sub-cavities. Nonetheless, this qualitative comparison hints that the condition for the transmissivity $T_{\text{c}}$ of the central mirror in Design B may resemble the condition for the reflectivity $R_{\text{c}}$ of the central mirror in Design A. Using an analogous derivation method as before, the following matrix equation for $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ can be established:
$\textbf{R}_{\text{CPA}}^{\text{(B)}}=\frac{t_{\text{a}}^{2}\,\textbf{T}_{\text{2f}}^{8}+(1+r_{\text{1a}}t_{\text{a}}^{2}-a)\,\textbf{T}_{\text{2f}}^{4}+r_{\text{1a}}\mathbb{1}}{r_{\text{1a}}t_{\text{a}}^{2}\,\textbf{T}_{\text{2f}}^{8}+(r_{\text{1a}}+t_{\text{a}}^{2}-a)\,\textbf{T}_{\text{2f}}^{4}+\mathbb{1}},$ (3)
with the scalar value $a=r_{\text{c}}^{2}(t_{\text{a}}^{2}+1)(r_{\text{1a}}^{2}+1)$. Observe how the structure of Eq. (3) resembles the structure of Eq. (1). Because $\textbf{T}_{\text{2f}}^{4}=\textbf{T}_{\text{4f}}^{2}$, the matrix $\textbf{T}_{\text{2f}}^{4}$ in Eq. (3) results in a quadruple Fourier transformation with a uniform phase shift $e^{2i\phi}$ for all modes, and can thus be represented as a simple diagonal matrix $\textbf{T}_{\text{4f}}^{2}=e^{2i\phi}\mathbb{1}$. Analogously, $\textbf{T}_{\text{2f}}^{8}=e^{4i\phi}\mathbb{1}$. Consequently, $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ becomes a diagonal matrix and can be reduced to a simple scalar equation, which proves the spatial degeneracy of Design B. The spectral degeneracy condition can be met when the following EP-conditions are fulfilled (see supplementary [45, S6]):
$R_{1\text{a}}$ $=T_{\text{a}}^{2}$ (4)
$R_{\mathrm{c}}=\left(\frac{1-T_{\mathrm{a}}}{1+T_{\mathrm{a}}}\right)^{2}\;\;$ $\Rightarrow\;\;T_{\mathrm{c}}=\frac{4T_{\mathrm{a}}}{(1+T_{\mathrm{a}})^{2}}$ (5)
Note the similarity between Eqs. (4), (5) and (2). While the central mirror in Design A must exhibit high reflectivity at the Exceptional Point (EP), in Design B it requires high transmissivity, which is in accordance with our earlier qualitative analysis. The analytical results confirm the anticipated quartic behavior and have been corroborated by numerical computer simulations, as illustrated in Fig. 4.
Conclusions. — We have introduced two CPA designs that combine spectral and spatial degeneracy. Through analytical and numerical models, we have validated the feasibility of this concept and explored its limitations. This new type of CPA is capable of perfectly depositing a light beam in a weak absorber over a wider spectral range as compared to a conventional CPA and regardless of the light beam’s spatial wavefront.
Acknowledgements — We would like to express our gratitude to Maximilian Prüfer (TU Wien), Jacob Scheuer (Tel Aviv University), and Birgitta Schultze-Bernhardt (TU Graz) for very helpful discussions. The numerical calculations were performed on the Vienna Scientific Cluster (VSC).